I am not sure exactly where to start on this homework problem. I would like you to please help me with this problem, so maybe I can get the rest of my homework.
Use log base 5 of 2=0.4307 and log base 5 of 3=0.6826 to approximate the value of each expression which is log base 5 of 9.
It is kind of difficult and if you understand what I am asking, please help me out on this.

Answer by mmm4444bot(95)

Hello There:
To write an expression for log_5(9), we change the 9 into 3^2. This allows us to use the property of logarithms which allows us to move the exponent into the coefficient position. In other words, the property is:
log(a^x) = x*log(a)
So, we have:
log_5(9) = log_5(3^2) = 2*log_5(3).
We're given a decimal approximation for log_5(3); double it to find an approximation for log_5(9).
~ Mark